# RL in Latent MDPs is Tractable: Online Guarantees via Off-Policy Evaluation

### v1

@@ 1.1 Technical Challenges
Many optimism based online RL algorithms follow a similar pattern. They make use of a confidence set – a set of candidate models (hypothesis) that can explain the observed data – and execute a policy that will shrink the volume of the confidence set is produced and executed [6, 37, 26, 7, 32, 35]. The entirety of the statistical problem is to analyze the decaying rate of confidence sets under proper model class assumptions [48, 33, 18].
@@
The work of [42] suggested an alternative strategy to learn a near-optimal policy in LMDPs: the moment-matching approach for exploration in LMDPs. When all contexts share the same state-transition dynamics, those moments can be defined as the joint distribution of rewards at a sequence of $d:=2M-1$ state-action pairs $\bm{x}=\left((s_{[1]},a_{[1]}),...,(s_{[d]},a_{[d]})\right)$. This in turn suggests that the exploration algorithm must learn how to visit these $d$ state-action pairs simultaneously, i.e., find a policy that ensures that $\bm{x}$ appears as a subsequence of the entire trajectory with high enough probability.
When the transition dynamics of different latent contexts is the similar, reaching optimally to $d$ state-action pairs minor challenge; e.g., we can first learn the shared transition kernel with any reward-free exploration scheme for MDPs [32], and then execute the policy that maximizes the probability of reaching to the $d$ state-action pairs. However, for general LMDP, when the transition dynamics of different latent contexts may differ, this approach is no longer. Furthermore, to follow the notion of moments suggested in [42], the data collected for estimating the correlation tensor must be collected under the same prior (belief) over all latent contexts. Unfortunately, ensuring this for general LMDPs, when the transition dynamics of different latent context is not equal, is impossible. Even if we obtain the samples of correlations, different policies may result in different priors over contexts. These challenges hint we may need an alternative approach to solve general LMDPs, when the transition dynamics vary between latent contexts.
@@ 1.2 Overview of Our Contribution
Recent studies have found fundamental connections between off-policy evaluation (OPE) and online exploration in RL [58, 2, 28, 9, 54, 3, 4]. In this work, we offer a fresh viewpoint, which deviates from existing works, on the connection between OPE and online exploration. This perspective, together with new analysis tools, allows us to provide a sample-efficient algorithm for the LMDP setting. This further showcases the usefulness of OPE for online exploration in POMDPs.
@@
How would this quantity be related to online exploration? A key observation towards start developing intuition is the following: an unbounded coverage coefficient, i.e., $C(\psi;\pi)=\infty$ implies there exists a state-action pair, at a some time-step, that cannot be reached under $\psi$, but reachable with $\pi$. However, for MDPs, the number of new state-action pairs is bounded, and with it, the number of times the coverage coefficient may be large during an interaction.
The algorithmic framework we lie in this work builds on OMLE [45]. In Section 3, we consider the MDP setting to provide intuition of our analysis. There, OMLE iteratively tests new policies on models from the confidence set which predict different outcomes, until the trajectory distribution of all policies is reliably estimated. We provide new analysis for the MDP setting based on OPE tools.
@@
where $m$ is the unobserved latent context, $\mathcal{T}=(s_{t},a_{t},r_{t})_{t\in[H]}$ is a sampled trajectory, $\mathcal{E}$ is an event of visiting a tuple of states and actions within an episode, and $\mathcal{I}$ is an intervention of interest, e.g., force an action $a$ at the $t^{th}$ time step regardless of $\psi$ (for the formal definition, see Definition 4.1).
Even though the coverage coefficient cannot be measured explicitly, since $m$ is a latent variable, it is central to our analysis and our ability to analyze the sample complexity of the proposed algorithm. Its usefulness lies in an OPE guarantee we develop (see Lemma 4.2):
$\texttt{TV}(\mathds{P}_{\theta^{*}}^{\pi},\mathds{P}_{\theta}^{\pi})(\mathcal{T})\lesssim C(\psi;\pi)\cdot\textstyle\sum_{\mathcal{I}}\texttt{TV}(\mathds{P}_{\theta^{*}}^{\psi},\mathds{P}_{\theta}^{\psi})(\mathcal{T}||\textbf{ do }\mathcal{I}),$
@@ 2 Preliminaries
Without loss of generality, we assume that there exists a null state that represents the starting and terminal state $s_{0}=s_{H+1}=\emptyset$, and a null action at the beginning of an episode $a_{0}=\emptyset$, even though actual policies do not take any action at the beginning. $T_{m}(\cdot|s_{0},a_{0})$ is the initial state distribution of the $m^{th}$ MDP. We assume that the number of latent contexts is constant $M=O(1)$, and the time-horizon is larger than the number of contexts $H>2M$ Further, we assume the reward values are finite and bounded:
@@
3 Online Exploration of MDPs via Off-Policy Evaluation: A New Perspective
In this section, we present our new approach for analyzing the OMLE algorithm, and, for establishing intuition in the Markovian setting. Differently than prior analysis [44, 45] which is based on the generalized eluder-type condition assumption (see [45], Condition 3.2), our approach utilizes a different structural property. Specifically, we show that the coverage coefficient and a certain type of an OPE guarantee can be used to study the performance of OMLE. This alternative perspective is instrumental in designing a sample-efficient algorithm for the LMDP class.
Consider MDP-OMLE depicted in Algorithm 1. MDP-OMLE is an adaptation of OMLE for the MDP setting with the goal of learning a near-optimal policy. The algorithm iteratively refines the confidence set, the set of statistically valid models, until halts. It iteratively repeats the two steps: (i) find a policy for which the TV distance between trajectory distributions of models in the confidence set is sufficiently large, and (ii) collect data with that policy and refine the confidence set. To bound the sample complexity of the algorithm we attempt to upper bound the number of iterations, namely, to bound the number of times the TV distance between trajectory distributions can be sufficiently large.
@@ 3 Online Exploration of MDPs via Off-Policy Evaluation: A New Perspective
The following OPE lemma is a new tool that allows us to bound the number of iterations of MDP-OMLE. Before discussing its application, we present the result.
@@
In a typical tabular MDP setting, we take $O(\log|\Theta|)=\tilde{O}(SA)$, by discretizing the class of MDPs. Hence the sample complexity of MDP-OMLE is $N=\tilde{O}(H^{6}S^{3}A^{3}/\epsilon^{2})$. While this upper bound is suboptimal compared to the minimax rate [7], the appeal of this type of analysis is its ability to bypass the need for shrinking confidence sets and the need to infer the belief state (Section 1.1, Challenge 1).
@@ 4.1 Off-Policy Evaluation in LMDPs
The OPE lemma we derive in this section makes use of a behavior policy of a special form which we refer as segmented policy. Let us formally define this construction. A segmented policy, which we denote by $\nu(\bm{\psi};{\bm{\tau}},\bm{z})$, takes as an input a sequence of history-dependent policies, $\bm{\psi}=(\psi_{0},...,\psi_{d})$, a sequence of time steps, we call checkpoints, $\bm{\tau}=(\tau_{1},...,\tau_{|\bm{\tau}|})\in\texttt{SubSeq}(H,d)$, and a sequence of binary numbers $\bm{z}=(z_{1},...,z_{|\bm{\tau}|})\in\{0,1\}^{|\bm{\tau}|}$ where $|\bm{\tau}|\leq d$, and returns a history-dependent policy.
@@
In Appendix C.4 we provide a counter-example where such single latent-state coverage coefficient is finite, and, yet, off-policy evaluation guarantee cannot be established.
@@ 4.2 Coverage Doubling via Sufficiency of Memoryless Polices
where $\nu(\psi_{i};\tau_{i})$ denotes a segmented policy executing $\psi_{i}$ after the $\tau_{i}^{th}$ time-step with memory reset, hence ignoring the history up to $\tau_{i}$ (the conditioning event $s_{\tau_{0}}^{\prime}=s_{[0]}^{\prime}$ at $i=0$ can be ignored). Inspired by the form in denominator, we may aim to double the following probability defined over a single segment:
@@ 4.3 The LMDP-OMLE Algorithm
    Find a memoryless policy $\pi^{k}\in\Pi_{\mathrm{mls}}$ whose prediction on trajectory distributions does not match between two models in the confidence set $\mathcal{C}^{k}$. Add $\pi^{k}$ to the collection of segment test policies, $\Psi^{k}_{\texttt{test}}$.
@@
    Collect new sample trajectories following the new set of segmented policies generated by different combination of collected test policies and switching operations.
@@
The data collection policy of LMDP-OMLE (second step above) is inspired and leverages Lemma 4.2 to give upper bounds on the TV distance of untested policies. Next, by Lemma 4.4, we know that when the while loop terminates, any optimal policy of a model contained in the confidence set is a near-optimal policy of the underlying LMDP.
We state our main theorem on the sample complexity of learning the optimal policy in Latent MDPs:
@@
Let $d=2M-1$ and assume $H>2M$. After at most $K=O(MS^{2}H)\cdot\log(MSAH/\epsilon)$ iterations, LMDP-OMLE terminates with at most $N$ episodes being generated where
@@ 5 Proof Sketch
    We cannot observe the pair of context and segment (the travel from one state to another) that previous behavioral policies could not cover.
For the first point, (a), we already have reduced the problem from matching all history-dependent policies to a set of behavioral policies generated by the concatenation of memoryless policies in Lemma 4.4. For the second point, (b), even though we cannot observe the latent context $m$ under which each segment is covered, we can improve the coverage of each segment given the test set $\Psi_{\texttt{test}}^{k}$ every iteration:
@@
for every $k^{th}$ iterations by generating sample trajectories with possible test policies from the set of segment policies. That is, we can ensure that the latent segment-coverage is being exponentially improved despite the unobservability of latent contexts $m$.
For a moment, to simplify the discussion, we first assume that the uniformly random policy $\texttt{Unif}(A)$ has a non-zero $\gamma>0$ segment probabilities in all contexts:
@@
We note that this assumption will be eventually removed in our final result. Thus, if we start from the segment-coverage $\gamma>0$, then since every probability is in the range of $[0,1]$, this doubling-up event can happen at most $\log(1/\gamma)$ times for every segment. Therefore, Algorithm 2 must terminate after at most $K=MS^{2}H\log(1/\gamma)$ iterations.
@@
Now we remove the $\gamma>0$ travel-coverage assumption (8). To do so, consider a virtual LMDP model $\hat{\theta}=(\{w_{m}^{*},\hat{T}_{m},R^{*}_{m}\})_{m=1}^{M}$ with perturbed transition models, i.e., $\hat{T}_{m}(\cdot|s,a)=(1-\gamma)T_{m}^{*}(\cdot|s,a)+\gamma\mathds{1}$ for all $(s,a)\in\mathcal{S}\times\mathcal{A}$. For $\hat{\theta}$, it is easy to see that for all $\pi\in\Pi$, we have $\texttt{TV}(\mathds{P}_{\hat{\theta}}^{\pi},\mathds{P}_{\theta^{*}}^{\pi})(\tau)\leq 2HS\gamma$. Thus, now we can shift our arguments to this perturbed model with sufficiently small $\gamma$, and it is straightforward that the perturbed model has $\gamma>0$ travel-coverage for any policy, which concludes Theorem 4.5.
@@ 6.1 Additional Details on the Full-Rankness Assumption
where $\psi_{\texttt{test}}\in\Psi_{\texttt{test}}$ is a test policy, and $\mathcal{T}_{t:H}$ is the future trajectory after time step $t$. Therefore, direct application of POMDP approaches such as OMLE require the prior knowledge of $\Psi_{\texttt{test}}$ and $\sigma_{\min}(L_{s})>0$ for all $s\in\mathcal{S}$. The rationale behind such assumptions is to ensure that a distribution of future trajectories can be converted to a belief of latent contexts, hence a distribution of future observations can serve as an alternative to a belief state. However, we are not given the set of core-tests $\Psi_{\texttt{test}}$, or even the existence of $\Psi_{\texttt{test}}$ that ensures $\sigma_{\min}(L_{s})>0$.

### v2

@@ 1.1 Technical Challenges
Many online RL algorithms follow a similar pattern. They make use of a confidence set – a set of candidate models (hypothesis) that can explain the observed data – and execute a policy that will shrink the volume of the confidence set is produced and executed [6, 38, 27, 7, 33, 36]. The entirety of the statistical problem is to analyze the decaying rate of confidence sets under proper model class assumptions [49, 34, 19].
@@
The work of [43] suggested an alternative strategy to learn a near-optimal policy in LMDPs: the moment-matching approach for exploration in LMDPs. When all contexts share the same state-transition dynamics, the notion of moments can be defined as the joint distribution of rewards under a fixed prior at a tuple of at most $d:=2M-1$ state-action pairs $\bm{x}=\left((s_{[1]},a_{[1]}),...,(s_{[d]},a_{[d]})\right)$. This in turn suggests that the exploration algorithm must learn how to visit these length-$d$ state-action tuples simultaneously, i.e., find a policy that ensures that $\bm{x}$ appears as a subsequence of the entire trajectory with high enough probability.
When the transition dynamics of different latent contexts is the similar, reaching optimally to $d$ state-action pairs is a minor challenge; e.g., we can first learn the shared transition kernel with any reward-free exploration scheme for MDPs [33], and then execute the policy that maximizes the probability of reaching to the $d$ state-action pairs. However, for general LMDP, when the transition dynamics of different latent contexts may differ, this approach is no longer available since the latent transition dynamics may not be learnable in general. Furthermore, to follow the notion of moments suggested in [43], the data collected for estimating the correlation tensor must be collected under the same prior (belief) over all latent contexts. Unfortunately, ensuring this for general LMDPs, when the transition dynamics of different latent context is not equal, is impossible, since even if we obtain the samples of correlations, different policies may result in different and unknown priors over contexts. These challenges hint we need an alternative approach to solve general LMDPs, when the transition dynamics vary between latent contexts.
@@ 1.2 Overview of Our Contribution
Recent studies have found some fundamental connections between off-policy evaluation (OPE) and online exploration in RL [59, 2, 29, 9, 55, 3, 4]. In this work, we offer a fresh viewpoint, which deviates from existing works, on the connection between OPE and online exploration. This perspective, together with new analysis tools, allows us to provide a sample-efficient algorithm for the LMDP setting. This further showcases the usefulness of OPE for online exploration in POMDPs.
@@
How would this quantity be related to online exploration? A key observation towards start developing intuition is the following: an unbounded coverage coefficient, i.e., $C(\psi;\pi)=\infty$ implies there exists a state-action pair, at some time-steps, that cannot be reached under $\psi$, but reachable with $\pi$.
The algorithmic framework we lie in this work builds on OMLE [46]. In Section 3, we consider the MDP setting to provide intuition of our analysis. There, OMLE iteratively tests new policies on models from the confidence set which predict different outcomes, until the trajectory distribution of all policies is reliably estimated. Since the number of new state-action pairs is bounded for MDPs, the number of times the coverage coefficient can be large must be bounded during an interaction. We provide new analysis for the MDP setting based on OPE tools.
@@
where $m$ is the unobserved latent context, $\mathcal{T}=(s_{t},a_{t},r_{t})_{t\in[H]}$ is a sampled trajectory, $\mathcal{E}$ is an event of visiting length of at most $d$ tuples of states and actions within an episode, and $\mathcal{I}$ is an intervention of interest, e.g., force an action $a$ at the $t^{th}$ time step regardless of $\psi$ (for the formal definition, see Definition 4.1). Note that the coverage coefficient cannot be measured explicitly, since $m$ is a latent variable; nevertheless, this concept is central to our analysis and our ability to analyze the sample complexity of the proposed algorithm. Its usefulness lies in an OPE guarantee we develop (see Lemma 4.2):
$\texttt{TV}(\mathds{P}_{\theta^{*}}^{\pi},\mathds{P}_{\theta}^{\pi})(\mathcal{T})\lesssim C(\psi;\pi)\cdot\textstyle\sum_{\mathcal{I}}\texttt{TV}(\mathds{P}_{\theta^{*}}^{\psi},\mathds{P}_{\theta}^{\psi})(\mathcal{T}\mid\textbf{do }\mathcal{I}),$
@@ 2 Preliminaries
Without loss of generality, we assume that there exists a null state that represents the starting and terminal state $s_{0}=s_{H+1}=\emptyset$, and a null action at the beginning of an episode $a_{0}=\emptyset$, even though actual policies do not take any action at the beginning. $T_{m}(\cdot|s_{0},a_{0})$ is the initial state distribution of the $m^{th}$ MDP. We assume that the number of latent contexts is constant $M=O(1)$, and the time-horizon is larger than the number of contexts $H>2M$. Further, we assume the reward values are finite and bounded:
@@
3 New Perspective on OMLE: Online Guarantees via Off-Policy Evaluation
In this section, we present our new approach for analyzing the OMLE algorithm, and, for establishing intuition in the Markovian setting. Differently than prior analysis [45, 46] which is based on the generalized eluder-type condition assumption (see [46], Condition 3.2), we utilize a certain type of an OPE guarantee can be used to study the performance of OMLE. This alternative perspective is instrumental in designing a sample-efficient algorithm for the LMDP class.
Consider MDP-OMLE depicted in Algorithm 1. MDP-OMLE is an adaptation of OMLE for the MDP setting with the goal of learning a near-optimal policy. The algorithm iteratively refines the confidence set, the set of statistically valid models, until it terminates. Specifically, it iteratively repeats the two steps: (i) find a policy for which the TV distance between trajectory distributions of models in the confidence set is sufficiently large, and (ii) collect data with that policy and refine the confidence set. To bound the sample complexity of the algorithm we attempt to upper bound the number of iterations, namely, to bound the number of times the TV distance between trajectory distributions can be sufficiently large.
@@ 3 New Perspective on OMLE: Online Guarantees via Off-Policy Evaluation
The following OPE lemma is a tool that allows us to bound the number of iterations of MDP-OMLE. Before discussing its application, we present the result.
@@
In a typical tabular MDP setting, we take $O(\log|\Theta|)=\tilde{O}(SA)$, by discretizing the class of MDPs. Hence the sample complexity of MDP-OMLE is $N=\tilde{O}(H^{6}S^{3}A^{3}/\epsilon^{2})$. While this upper bound is suboptimal compared to the minimax rate [7], the appeal of this type of analysis is its ability to bypass the need for analyzing the decaying rate of confidence sets and the need to infer the belief state (Section 1.1, Challenge 1).
@@ 4 Exploration in LMDPs via Sufficient Coverage
Intuition from moment-exploration algorithm in [43].
Before we dive into our key results, let us provide our intuition on how we construct the OPE lemma for LMDPs. Our construction is inspired by the moment-exploration algorithm proposed in [43]: when state-transition dynamics are identical across latent contexts, i.e., $T_{1}=T_{2}=...=T_{M}$, we can first learn the transition dynamics with any reward-free type exploration scheme for MDPs [33], and then set the exploration policy that sufficiently visits some tuples of state-actions $\bm{x}$ of length at most $d$. Specifically, they set a memorlyess exploration policy $\psi\in\Pi_{\texttt{mls}}$ which sets $\mathds{P}^{\psi}(x_{\bm{\tau}}=\bm{x})$ sufficiently large for some $\bm{\tau}\in\texttt{SubSeq}(H,d)$ and $\bm{x}\in\mathcal{X}^{\bigotimes|\bm{\tau}|}$. We note that the same moment-exploration strategy cannot be applied to general LMDPs with different state-transition dynamics since learning the transition dynamics itself involves latent contexts. Nevertheless, the intuition from [43] suggests that our key statistics are this visitation probabilities to all tuples of state-actions within a trajectory.
@@ 4.1 Off-Policy Evaluation in LMDPs
The OPE lemma we derive in this section makes use of a behavior policy of a special form which we refer as segmented policy, inspired by the notion of moment-exploration in [43]. Let us formally define the key quantities to establish our OPE lemma. A segmented policy, which we denote by $\nu(\bm{\psi};{\bm{\tau}},\bm{z})$, takes as an input a sequence of history-dependent policies, $\bm{\psi}=(\psi_{0},...,\psi_{d})$, a sequence of time steps, we call checkpoints, $\bm{\tau}=(\tau_{1},...,\tau_{|\bm{\tau}|})\in\texttt{SubSeq}(H,d)$, and a sequence of binary numbers $\bm{z}=(z_{1},...,z_{|\bm{\tau}|})\in\{0,1\}^{|\bm{\tau}|}$ where $|\bm{\tau}|\leq d$, and returns a history-dependent policy.
@@
In Appendix C.4 we provide a counter-example where such single latent-state coverage coefficient is finite, and yet, off-policy evaluation guarantee cannot be established.
@@ 4.2 Coverage Doubling via Sufficiency of Memoryless Polices
where $\nu(\psi_{i};\tau_{i})$ denotes a segmented policy executing $\psi_{i}$ after the $\tau_{i}^{th}$ time-step with memory reset, hence ignoring the history up to $\tau_{i}$ (the conditioning event $s_{\tau_{0}}^{\prime}=s_{[0]}^{\prime}$ at $i=0$ can be ignored). Inspired by the form in denominator, we aim to double the following probability defined over a context-segment pair:
@@ 4.3 The LMDP-OMLE Algorithm
    Find a memoryless policy $\pi^{k}\in\Pi_{\mathrm{mls}}$ whose prediction on trajectory distributions does not match between two models in the confidence set $\mathcal{C}^{k}$. Add $\pi^{k}$ to the collection of test policies $\Psi^{k}_{\texttt{test}}$, that forms each segment of (segmented) exploration policies.
@@
    Collect new sample trajectories following the new set of segmented policies for exploration, generated by different combination of collected test policies and switching operations.
@@
The data collection policy of LMDP-OMLE (second step above) is inspired and leverages Lemma 4.2 to give upper bounds on the TV distance of untested policies. Next, by Lemma 4.4, we know that when the while loop terminates, any optimal policy of a model contained in the confidence set is a near-optimal policy of the underlying LMDP. We conclude this section with our main theorem on the sample complexity of learning the optimal policy in Latent MDPs:
@@
Let $d=2M-1$ and assume $H>2M$. After at most $K=O(MS^{2}H)\cdot\log(MSAH/\epsilon)$ iterations, LMDP-OMLE (Algorithm 2) terminates with at most $N$ episodes being generated where
@@ 5 Proof Sketch
    We cannot observe any context-segment pair that previous behavioral policies could not cover.
For the first point, (a), we already have reduced the problem from matching all history-dependent policies to a set of behavioral policies generated by the concatenation of memoryless policies in Lemma 4.4. For the second point, (b), even though we cannot observe the latent context $m$ under which each segment is covered, we can improve the coverage of each context-segment pair given the test set $\Psi_{\texttt{test}}^{k}$ every iteration:
@@
That is, we can ensure that the coverage of at least one context-segment pair is being exponentially improved despite the unobservability of latent contexts $m$.
For a moment, to simplify the discussion, we first assume that the uniformly random policy $\texttt{Unif}(A)$ has a non-zero $\gamma>0$ probability for covering all segments in all contexts:
@@
We note that this assumption will be eventually removed in our final result. Thus, if we start from the initial coverage $\gamma>0$ over all context-segment pairs, then since every probability is in the range of $[0,1]$, this doubling-up event can happen at most $\log(1/\gamma)$ times for every context-segment pair. Therefore, Algorithm 2 must terminate after at most $K=MS^{2}H\log(1/\gamma)$ iterations.
@@
Now we remove the initial $\gamma>0$ coverage assumption (8). To do so, consider a virtual LMDP model $\hat{\theta}=(\{w_{m}^{*},\hat{T}_{m},R^{*}_{m}\})_{m=1}^{M}$ with perturbed transition models, i.e., $\hat{T}_{m}(\cdot|s,a)=(1-\gamma)T_{m}^{*}(\cdot|s,a)+\gamma\mathds{1}$ for all $(s,a)\in\mathcal{S}\times\mathcal{A}$. For $\hat{\theta}$, it is easy to see that for all $\pi\in\Pi$, we have $\texttt{TV}(\mathds{P}_{\hat{\theta}}^{\pi},\mathds{P}_{\theta^{*}}^{\pi})(\tau)\leq 2HS\gamma$. Thus, now we can shift our arguments to this perturbed model with sufficiently small $\gamma$, and it is straightforward that the perturbed model has the $\gamma>0$ segment coverage for any policy, which concludes Theorem 4.5.
@@ 6.1 Additional Details on the Full-Rankness Assumption
where $\psi_{\texttt{test}}\in\Psi_{\texttt{test}}$ is a test policy, and $\mathcal{T}_{t:H}$ is the future trajectory after time step $t$. Therefore, direct application of POMDP approaches such as OMLE require the prior knowledge of $\Psi_{\texttt{test}}$ and $\sigma_{\min}(L_{s})>0$ for all $s\in\mathcal{S}$. The rationale behind such assumptions is to ensure that a distribution of future trajectories can be converted to a belief of latent contexts, hence a distribution of future observations can serve as an alternative to a belief state. However, we are not given the set of core-tests $\Psi_{\texttt{test}}$, or even the existence of $\Psi_{\texttt{test}}$ that ensures $\sigma_{\min}(L_{s})>0$ for general Latent MDPs.
With Separation.
In a recent work by Chen et al. [14], a polynomial upper bound in $M$ has been established under a notion of strong-separation assumption between contexts with a sufficiently long time horizon $H$. In essence, their assumptions guarantee that $\sigma_{\min}(L_{s})>0$ holds for most of the states with a priori given test policy, along with additional analysis for the tail of episodes. It is of great importance to identify such practical assumptions that lead to fully polynomial upper bounds, especially for instances with some proper notion of separations even when no prior knowledge of test policies is provided.